### v2

@@ 3.2 Dirac-GAN problem
Results. Figure 2 shows the stationarity gap and the sequences generated by all three test algorithms. We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point. Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of the AGP algorithm.
@@ 3.3 Robust learning over multiple domains
where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$. We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky et al., 2009). Our goal is to train a neural network that works on these two completely unrelated problems simultaneously. Since cross-entropy loss is popular in multi-class classification problem, we use it as our loss function. We should point out that the quality of the algorithm for solving a robust learning problem over multiple domains is measured by the worst case accuracy over all domains (Qian et al., 2019).
Experiment setup. We compare the proposed PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$, $l_{12}^{1}=0.01$ and $l_{22}^{1}=0.01$. For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20. For the AGP algorithm, we set $\frac{1}{\beta_{k}}=\frac{2}{2+\sqrt{k}}$, $\gamma_{k}=100$ and $c_{k}=\frac{1}{10+k^{\frac{1}{4}}}$. We set the batch size in all tasks to be $128$ and run $50$ epochs for all algorithms. Moreover, we sample our results every epoch and calculate the average accuracy on these two testing sets to evaluate the performance of different algorithms. Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
Figure 3: Performance of two algorithms for solving robust multi-task learning problem.
Results. Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively. PF-AGP-NC algorithm outperforms the SGDA-B algorithm and the AGP algorithm for both datasets.
@@ 4 Conclusions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results demonstrate the efficiency of the two proposed algorithms.

### v3

@@ 4.2 Dirac-GAN problem
Results. Figure 2 shows the stationarity gap and the sequences generated by all three test algorithms. We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point. Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of AGP algorithm.
@@ 4.3 Robust learning over multiple domains
where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$. We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky and Hinton, 2009). The dimension of $x$ is 57,044,810. The sample sizes of the CIFAR-10 and MNIST datasets are both 60,000. Our goal is to train a neural network that works on these two completely unrelated problems simultaneously. Since cross-entropy loss is popular in multi-class classification problem, we use it as our loss function. We should point out that the quality of the algorithm for solving a robust learning problem over multiple domains is measured by the worst case accuracy over all domains (Qian et al., 2019).
Experiment setup. We compare the proposed PF-AGP-NC algorithm, PF-AGP-NL algorithm with the SGDA-B algorithm and the AGP algorithm. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$ and $l_{22}^{1}=5$. For the PF-AGP-NL algorithm, we set $l_{11}^{1}=0.1$ and $l_{12}^{1}=0.1$. For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20. For the AGP algorithm, we set $\frac{1}{\beta_{k}}=\frac{2}{2+\sqrt{k}}$, $\gamma_{k}=100$ and $c_{k}=\frac{1}{10+k^{\frac{1}{4}}}$. We set the batch size in all tasks to be $128$ and run $50$ epochs for all algorithms. Moreover, we sample our results every epoch and calculate the average accuracy on these two testing sets to evaluate the performance of different algorithms. Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
Figure 3: Numerical Performance of four tested algorithms for solving robust multi-task learning problem.
Results. Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively. The PF-AGP-NL algorithm and PF-AGP-NC algorithm outperform the SGDA-B algorithm and the AGP algorithm for both datasets.
@@ 5 Conclusions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concavity modulus $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by  $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$.
Moreover, we also propose a completely parameter-free alternating gradient projection algorithm, i.e., the PF-AGP-NL algorithm, for solving nonconvex-linear minimax problems, which does not require any prior knowledge of parameters. As far as we know, the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm and the PF-AGP-NL algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems, nonconvex-concave minimax problems and nonconvex-linear minimax problems respectively. Numerical results demonstrate the efficiency of the proposed algorithms.
Appendix A A restarted parameter-free single-loop algorithm for nonconvex-strongly concave minimax problems
In this section, we propose a restarting parameter-free alternating gradient projection algorithm (rPF-AGP-NSC) for the nonconvex-strongly concave minimax problem. The algorithm is a restarting variant of PF-AGP-NSC that does not require the prior knowledge of the Lipschtiz constant $L$, but requires the prior knowledge of two additional parameters $\mu$ and $\underline{S}$. The algorithm is formally presented in Algorithm 4.
Next, we prove the iteration complexity of Algorithm 4 under the nonconvex-strongly concave setting. Denote $\Phi(x):=\max_{y\in\mathcal{Y}}f(x,y)$ and $y^{*}(x):=\arg\max_{y\in\mathcal{Y}}f(x,y)$. By Lemma B.1 in (Nouiehed et al., 2019a), $\Phi(x)$ is $L_{\Phi}$-Lipschitz smooth with $L_{\Phi}=L+\frac{L^{2}}{\mu}$. Moreover, we have
$\nabla_{x}\Phi(x)=\nabla_{x}f(x,y^{*}(x)).$ (120)
Algorithm 4 A restarted parameter-free alternating gradient projection (rPF-AGP-NSC) algorithm for nonconvex-strongly concave minimax problems
  Step 1: Input $x_{1,1},y_{1,1}$, $l_{1}$, $\mu$, $\varepsilon$, $\underline{S}$; Set $i=1$. Compute: $S_{1}=2\max_{y\in\mathcal{Y}}f(x_{1,1},y)-f(x_{1,1},y_{1,1})$.
  Step 2: Comppute $\beta_{i}=3l_{i}$, $\alpha_{i}=\frac{158l_{i}^{3}}{\mu^{2}}$, $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+l_{i})}{\mu\beta_{i}}$, $\tilde{d}^{i}_{1}=\min\{\frac{\alpha_{i}-\frac{l_{i}(l_{i}+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{l_{i}^{2}}{\mu}-\frac{3l_{i}}{2}}{2\alpha_{i}^{2}},\frac{\beta_{i}-\frac{3l_{i}}{2}}{\beta_{i}^{2}+2l_{i}^{2}}\}$, $\tilde{T}_{i}=\frac{S_{1}-\underline{S}}{\varepsilon^{2}\tilde{d}^{i}_{1}}$; Set $k=1$;
  Step 3: Update $x_{i,k+1}$ and $y_{i,k+1}$:
   (a): If $k\geq\tilde{T}_{i}$ then, go to Step 4;
   (b): Update $x_{i,{k+1}}$ and $y_{i,{k+1}}$
$y_{i,{k+1}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{i,k}+\frac{1}{\beta_{i}}\nabla_{y}f(x_{i,{k}},y_{i,k})\right),$ (121)
$x_{i,{k+1}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{i,k}-\frac{1}{\alpha_{i}}\nabla_{x}f(x_{i,k},y_{i,k+1})\right).$ (122)
   (c): Set $k=k+1$ go to Step 3.
  Step 4: If $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$, stop; Otherwise $x_{i+1,1}=x_{1,1}$, $y_{i+1,1}=y_{1,1}$, $l_{i+1}=2l_{i}$, $i=i+1$, go to Step 2.
Lemma 17
Suppose that Assumption 1 holds. Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Let $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+L)}{\mu\beta_{i}}$, then $\forall k,i\geq 1$, and for any $a>0$,
$\Phi(x_{i,k+1})-\Phi(x_{i,k})$
$\leq$ $\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle+\frac{L^{2}(L+\beta_{i})^{2}\eta_{i}^{2}}{2a\beta_{i}^{2}}\|y_{i,k+1}-y_{i,k}\|^{2}+\frac{a+L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (123)
Proof  Since $\Phi(x)$ is $L_{\Phi}$ smooth with respect to $x$, and by (120), we have
$\Phi(x_{i,k+1})-\Phi(x_{i,k})\leq$ $\langle\nabla_{x}f(x_{i,k},y^{*}(x_{i,k}))-\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle$
$+\frac{L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle.$ (124)
By the Cauchy-Schwarz inequality, for any $a>0$, we obtain
$\langle\nabla_{x}f(x_{i,k},y^{*}(x_{i,k}))-\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle$
$\leq$ $\frac{L^{2}}{2a}\|y_{i,k+1}-y^{*}(x_{i,k})\|^{2}+\frac{a}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (125)
By Lemma 1 in (Zhang et al., 2022), we have
$\|y_{i,k+1}-y^{*}(x_{i,k})\|\leq\eta_{i}\left\|y_{i,k+1}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k+1}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k+1}\right)\right)\right\|.$ (126)
By (121), the non-expansive property of the projection operator $\mathcal{P}_{\mathcal{Y}}(\cdot)$ and Assumption 1, we obtain
$\left\|y_{i,k+1}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k+1}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k+1}\right)\right)\right\|\leq\frac{L+\beta_{i}}{\beta_{i}}\|y_{i,k+1}-y_{i,k}\|.$ (127)
Combing (A), (125), (126) and (127), we get
$\Phi(x_{i,k+1})-\Phi(x_{i,k})\leq$ $\frac{L^{2}\eta_{i}^{2}(L+\beta_{i})^{2}}{2a\beta_{i}^{2}}\|y_{i,k+1}-y_{i,k}\|^{2}+\frac{a+L_{\Phi}}{2}\|x_{i,k+1}-x_{i,k}\|^{2}$
$+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle,$ (128)
which completes the proof.
Next, we provide an lower bound for the difference between $f(x_{i,k+1},y_{i,k+1})$ and $f(x_{i,k},y_{i,k})$.
Lemma 1
Suppose that Assumption 1 holds. Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Then $\forall i,k\geq 1$,
$f(x_{i,k+1},y_{i,k+1})-f(x_{i,k},y_{i,k})$
$\geq$ $\left(\beta_{i}-\frac{L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}-\frac{L}{2}\|x_{i,k+1}-x_{i,k}\|^{2}+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle,$ (129)
Proof  The optimality condition for $y_{i,k}$ in (121) implies that $\forall y\in\mathcal{Y}$ and $\forall i,k\geq 1$,
$\langle\nabla_{y}f(x_{i,k},y_{i,k}),y_{i,k+1}-y_{i,k}\rangle\geq\beta_{i}\|y_{i,k+1}-y_{i,k}\|^{2}.$ (130)
By Assumption 1, the gradient of $f(x,y)$ is Lipschitz continuous with respect to $y$ and (130), we get
$f(x_{i,k},y_{i,k+1})-f(x_{i,k},y_{i,k})\geq$ $\left(\beta_{i}-\frac{L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}.$ (131)
Since that the gradient of $f(x,y)$ is Lipschitz continuous with respect to $x$, we obtain
$f(x_{i,k+1},y_{i,k+1})-f(x_{i,k},y_{i,k+1})$
$\geq$ $\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle-\frac{L}{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (132)
The proof is then completed by adding (131) and (A).
We now establish an important recursion for the Algorithm 4.
Lemma 2
Suppose that Assumption 1 holds. Denote $S(x,y)=2\Phi(x)-f(x,y).$ Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Then $\forall i,k\geq 1$,
$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})\leq$ $-\left(\alpha_{i}-\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{L^{2}}{\mu}-\frac{3L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}$
$-\left(\beta_{i}-\frac{3L}{2}\right)\|y_{i,k+1}-y_{i,k}\|^{2}.$ (133)
Proof  Denote $S(x,y)=2\Phi(x)-f(x,y)$. Combining (17) and (1), we have
$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})$
$\leq$ $-\left(\beta_{i}-\frac{L}{2}-\frac{(L+\beta_{i})^{2}\eta_{i}^{2}L^{2}}{a\beta_{i}^{2}}\right)\|y_{i,k+1}-y_{i,k}\|^{2}+\left(a+L_{\Phi}+\frac{L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}$
$+\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle.$ (134)
The optimality condition for $x_{i,k}$ in (122) implies that $\forall x\in\mathcal{X}$ and $\forall k\geq 1$,
$\langle\nabla_{x}f(x_{i,k},y_{i,k+1}),x_{i,k+1}-x_{i,k}\rangle\leq-\alpha_{i}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (135)
Plugging (135) into (A), then we obtain
$S(x_{i,k+1},y_{i,k+1})-S(x_{i,k},y_{i,k})$
$\leq$ $-\left(\beta_{i}-\frac{L}{2}-\frac{L^{2}(L+\beta_{i})^{2}\eta^{2}}{a\beta_{i}^{2}}\right)\|y_{i,k+1}-y_{i,k}\|^{2}-\left(\alpha_{i}-a-L_{\Phi}-\frac{L}{2}\right)\|x_{i,k+1}-x_{i,k}\|^{2}.$ (136)
The proof is then completed by the definiton of $L_{\Phi}=L+\frac{L^{2}}{\mu}$ and choosing $a=\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}$ in (A).
We are now ready to establish the iteration complexity for the rPF-AGP-NSC algorithm. In particular, letting $\nabla G(x_{i,k},y_{i,k})$ be defined as in Definition 1 and $\varepsilon>0$ be a given target accuracy, we provide a bound on $T_{i}(\varepsilon)$, the inner iteration index to achieve an $\varepsilon$-stationary point, i.e., $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$, which is equivalent to
$T_{i}(\varepsilon)=\min\{k\mid\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon\}.$ (137)
The following theorem gives an upper bound on $T_{i}(\varepsilon)$.
Theorem 1
Suppose that Assumptions 1 hold. Let $\{\left(x_{i,k},y_{i,k}\right)\}$ be a sequence generated by Algorithm 4. Let $\eta_{i}=\frac{(2\beta_{i}+\mu)(\beta_{i}+L)}{\mu\beta_{i}}$. If
$\beta_{i}>\frac{3L}{2},\alpha_{i}>\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}+\frac{L^{2}}{\mu}+\frac{3L}{2},$ (138)
then $\forall\varepsilon>0$, we have
$T_{i}\left(\varepsilon\right)\leq\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}},$
where $d^{i}_{1}:=\min\{{\frac{\alpha_{i}-\frac{L(L+\beta_{i})^{2}\eta_{i}^{2}}{\beta_{i}^{2}}-\frac{L^{2}}{\mu}-\frac{3L}{2}}{2\alpha_{i}^{2}},\frac{\beta_{i}-\frac{3L}{2}}{\beta_{i}^{2}+2L^{2}}\}}$, $d_{2}^{i}:=S(x_{i,1},y_{i,1})-\underline{S}$ with $\underline{S}:=\min\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}f(x,y)$.
Proof  By (121), we immediately obtain
$\left\|\beta_{i}\left(y_{i,k}-\mathcal{P}_{\mathcal{Y}}\left(y_{i,k}+\frac{1}{\beta_{i}}\nabla_{y}f\left(x_{i,k},y_{i,k},\right)\right)\right)\right\|=\beta_{i}\|y_{i,k+1}-y_{i,k}\|.$ (139)
On the other hand, by (122) and the Cauchy-Schwartz inequality, we have
$\|\alpha_{i}(x_{i,k}-\mathcal{P}_{\mathcal{Y}}(x_{i,k}-\frac{1}{\alpha_{i}}\nabla_{x}f\left(x_{i,k},y_{i,k}\right)))\|\leq\alpha_{i}\|x_{i,k+1}-x_{i,k}\|+L\|y_{i,k+1}-y_{i,k}\|,$ (140)
where the last inequality is by the nonexpansive property of the projection operator and Assumption 1. Combing (139), (140), and using Cauchy-Schwarz inequality, we have
$\|\nabla G(x_{i,k},y_{i,k})\|^{2}\leq(\beta_{i}^{2}+2L^{2})\|y_{i,k+1}-y_{i,k}\|^{2}+2\alpha_{i}^{2}\|x_{i,k+1}-x_{i,k}\|^{2}.$ (141)
By (138), it can be easily checked that $d^{i}_{1}>0$. By multiplying $d^{i}_{1}$ on the both sides of (141) and using (2) in Lemma 2, we get
$d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}\leq S(x_{i,k},y_{i,k})-S(x_{i,k+1},y_{i,k+1}).$ (142)
Summing both sides of (142) from $k=1$ to $T_{i}(\varepsilon)$, we then obtain
$\sum_{k=1}^{T_{i}\left(\varepsilon\right)}{d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}}\leq S(x_{i,1},y_{i,1})-S(x_{i,T_{i}(\varepsilon)+1},y_{i,T_{i}(\varepsilon)+1}).$ (143)
Then, by (143) we obtain $\sum_{k=1}^{T_{i}\left(\varepsilon\right)}{d^{i}_{1}\|\nabla G(x_{i,k},y_{i,k})\|^{2}}\leq S(x_{i,1},y_{i,1})-\underline{S}=d_{2}^{i}.$ In view of the definition of $T_{i}(\varepsilon)$, the above inequality implies that $\varepsilon^{2}\leq\frac{d_{2}^{i}}{T_{i}(\varepsilon)d^{i}_{1}}$ or equivalently, $T_{i}\left(\varepsilon\right)\leq\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}}$. This completes the proof.
Remark 18
Let $\beta_{i}=3l_{i},\alpha_{i}=\frac{158l_{i}^{3}}{\mu^{2}}$, then the outer iteration of the rPF-AGP-NSC algorithm requires at most $\mathcal{O}(\log(L))$ steps, so that $\beta_{i}$ and $\alpha_{i}$ satisfy (138). According to Theorem 1, the inner iteration of $\|\nabla G(x_{i,k},y_{i,k})\|\leq\varepsilon$ requires at most $\frac{d_{2}^{i}}{\varepsilon^{2}d^{i}_{1}}$ steps.
Remark 19
Denote $\kappa=L/\mu$. If we set $\beta_{i}=3l_{i}$, by Theorem 1 we have that $\eta_{i}=\mathcal{O}(\kappa)$, $d_{1}^{i}=\mathcal{O}(\frac{1}{L\kappa^{2}})$ and $d_{2}^{i}=O(1)$, which implies the total number of gradient calls is upper bounded by $\sum\limits_{i=1}^{\log_{2}(L)+1}T_{i}(\varepsilon)=\mathcal{O}(\log(L)L\kappa^{2}\varepsilon^{-2})$ under the nonconvex-strongly concave setting. It means that if two additional parameters $\mu$ and $\underline{S}$ are known in advance, the gradient complexity of the rPF-AGP-NSC algorithm is improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than that of the SGDA-B and the PF-AGP-NSC algorithm.